# CARTE: pretraining and transfer for tabular learning

### v1

@@ B.3 Datasets from multi-table experiments
Wine prices
: Wina Poland, WineEnthusiasts, WineVivino, Wine.com
Wine ratings
: WineEnthusiasts, WineVivino, Wine.com
Beers
: Beer Ratings, Wikiliq-Beer
Used Cars
Used Cars 24, Used Cars Benz Italy, UsedCars.com, Used Cars Pakistan, Used Cars Saudi Arabia
Films
Filmtv Movies, Rotten Tomatoes
Dramas
K-Drama, Mydramalist
Animes
Anime Planet, Japanese Anime
Baby products
Buy Buy Baby, Babies R Us
Bike sales
Bikedekho, Bikewale
Employee remunerations
Company Employees, Employee remuneration and expenses earning over 75000, ML/DS Salaries
Restaurant ratings
Zomato, Michelin, Yelp
Journal scores
Journal Score JCR, Journal Score SJR
@@ Appendix C Extended results
C.1 The effect of oversmoothing
@@
Figure 8 show that representations extracted from deeper layers of the Graph Neural Networks (GNNs) are less useful for prediction on downstream tasks [^46] Here they are used inside a HistGradientBoosting predictor from scikit-learn.. We interpret this as an effect of oversmoothing, a well known problem in GNNs (Chen et al., 2020a; Rusch et al., 2023).
@@
C.2 Details of the entity matching experiment
@@
C.3 Schema-matching results on joint learning across multiple tables
Figure 9 depicts a direct comparison of performance between CARTE with and without schema-matching over 275 different cases in number of source data (ranging from one source to five sources). Each point represents a comparison of the average score over a dataset with different train/test split for a given size of train data. If a point is located below the diagonal line, it indicates a higher performance of x-axis. From the figure, we see that dots align along the diagonal line, indicating similar performance between CARTE with and without schema-matching (also with p-value of 0.728 for two-sided t-test on difference in means). The results suggest schema-matching is not required for CARTE on transfer across multiple tables.

### v2

@@ B.3 Datasets from multi-table experiments
Wine prices: Wina Poland, WineEnthusiasts, WineVivino, Wine.com
Wine ratings: WineEnthusiasts, WineVivino, Wine.com
Beers: Beer Ratings, Wikiliq-Beer
Used Car: Used Cars 24, Used Cars Benz Italy, UsedCars.com, Used Cars Pakistan, Used Cars Saudi Arabia
Films: Filmtv Movies, Rotten Tomatoes
Dramas: K-Drama, Mydramalist
Animes: Anime Planet, Japanese Anime
Baby products: Buy Buy Baby, Babies R Us
Bike sales: Bikedekho, Bikewale
Employee remunerations: Company Employees, Employee remuneration and expenses earning over 75000, ML/DS Salaries
Restaurant ratings: Zomato, Michelin, Yelp
Journal scores: Journal Score JCR, Journal Score SJR
@@ Appendix C Extended Results
C.1 Performance Comparison of CARTE with 42 Baselines for Learning on Single Tables
As an extension to the results shown in subsection 4.2, Figure 9 shows the overall comparison of CARTE with 42 baselines that additionally accounts for Scikit-Learn’s HistGradientBoosting (HGB), target encoding for categorical variables (Micci-Barreca, 2001) (TarEnc), employing external information from language models of Fasttext (FT) and intfloat/e5-small-v2 (Wang et al., 2022) (LLM), and the ‘Bagging’ strategy. We see that CARTE attains the pronounced lead to all the baselines for both regression and classification tasks. Moreover, it is interesting to observe that the bagging strategy has positive impacts for neural network models, while the effect is limited on linear or ensembling baselines (tree-based models or TabPFN). This may hint that bagging with different train/validation splits is an important setups for other deep learning architectures, especially for limited train-sizes.
(a) Regression
(b) Classification
Figure 9: Comparison of CARTE with 42 baselines on single tables. The critical difference diagram for CARTE and 42 baselines for (a) regression (b) classification tasks. In addition to the methods in Figure 4 (with the same coloring), we include additional baselines with HistGradientBoosting, target encoding, and external information from language models. The figure shows that CARTE attains the pronounced lead to all the baselines for both regression and classification tasks. Moreover, the bagging strategy brings larger benefits to neural networks compared to linear or ensembling baselines, which suggests an important setups for other deep learning architectures.
Table 5: Detailed results of TabLLM datasets. Dataset specification and performance comparison among CARTE and three baselines presented in TabLLM (Hegselmann et al., 2023). The datasets contain high fraction of numerical features or categorical columns with low cardinality. As observed from the comparison results, TabPFN performs the best followed by CARTE, XGBoost, and TabLLM.
TabLLM dataset specificiations
Datasets Fraction of Average
numerical columns cardinality
bank 0.40 38.44
blood 1.00 0.00
calhousing 1.00 0.00
car 0.00 3.50
creditg 0.33 3.79
diabetes 1.00 0.00
heart 0.45 2.67
income 0.33 12.38
jungle 1.00 0.00
Performance comparison
Datasets Methods Number of Shots
32 64 128 256 512
bank CARTE 0.81$\pm$0.03 0.83$\pm$0.03 0.87$\pm$0.04 0.89$\pm$0.03 0.90$\pm$0.01
XGBoost 0.76$\pm$0.03 0.83$\pm$0.02 0.85$\pm$0.03 0.88$\pm$0.01 0.90$\pm$0.01
TabPFN 0.76$\pm$0.03 0.82$\pm$0.03 0.86$\pm$0.02 0.89$\pm$0.00 0.90$\pm$0.00
TabLLM 0.64$\pm$0.06 0.69$\pm$0.03 0.82$\pm$0.05 0.87$\pm$0.01 0.88$\pm$0.01
blood CARTE 0.68$\pm$0.01 0.68$\pm$0.01 0.72$\pm$0.02 0.72$\pm$0.0 0.71$\pm$0.01
XGBoost 0.67$\pm$0.06 0.68$\pm$0.05 0.71$\pm$0.06 0.70$\pm$0.07 0.67$\pm$0.06
TabPFN 0.70$\pm$0.04 0.73$\pm$0.04 0.75$\pm$0.04 0.76$\pm$0.04 0.76$\pm$0.03
TabLLM 0.68$\pm$0.04 0.68$\pm$0.04 0.68$\pm$0.06 0.70$\pm$0.08 0.68$\pm$0.04
calhousing CARTE 0.79$\pm$0.02 0.83$\pm$0.03 0.85$\pm$0.04 0.87$\pm$0.05 0.89$\pm$0.05
XGBoost 0.79$\pm$0.04 0.82$\pm$0.04 0.87$\pm$0.01 0.90$\pm$0.01 0.92$\pm$0.01
TabPFN 0.85$\pm$0.03 0.89$\pm$0.01 0.91$\pm$0.01 0.92$\pm$0.00 0.93$\pm$0.00
TabLLM 0.77$\pm$0.08 0.77$\pm$0.04 0.81$\pm$0.02 0.83$\pm$0.01 0.86$\pm$0.02
car CARTE 0.87$\pm$0.06 0.94$\pm$0.07 0.98$\pm$0.03 0.99$\pm$0.03 1.00$\pm$0.02
XGBoost 0.82$\pm$0.03 0.91$\pm$0.02 0.95$\pm$0.01 0.98$\pm$0.01 0.99$\pm$0.01
TabPFN 0.92$\pm$0.02 0.97$\pm$0.00 0.99$\pm$0.01 1.00$\pm$0.00 1.00$\pm$0.00
TabLLM 0.91$\pm$0.02 0.96$\pm$0.02 0.98$\pm$0.01 0.99$\pm$0.00 1.00$\pm$0.00
creditg CARTE 0.67$\pm$0.03 0.68$\pm$0.01 0.70$\pm$0.02 0.75$\pm$0.01 0.77$\pm$0.02
XGBoost 0.66$\pm$0.03 0.67$\pm$0.06 0.68$\pm$0.02 0.73$\pm$0.02 0.75$\pm$0.03
TabPFN 0.69$\pm$0.07 0.70$\pm$0.07 0.72$\pm$0.06 0.75$\pm$0.04 0.75$\pm$0.02
TabLLM 0.72$\pm$0.06 0.70$\pm$0.07 0.71$\pm$0.07 0.72$\pm$0.03 0.72$\pm$0.02
diabetes CARTE 0.76$\pm$0.06 0.79$\pm$0.02 0.81$\pm$0.01 0.82$\pm$0.01 0.81$\pm$0.00
XGBoost 0.69$\pm$0.08 0.73$\pm$0.05 0.78$\pm$0.05 0.80$\pm$0.03 0.80$\pm$0.01
TabPFN 0.77$\pm$0.03 0.82$\pm$0.03 0.83$\pm$0.03 0.83$\pm$0.03 0.81$\pm$0.02
TabLLM 0.68$\pm$0.04 0.73$\pm$0.03 0.79$\pm$0.04 0.78$\pm$0.02 0.78$\pm$0.04
heart CARTE 0.90$\pm$0.02 0.91$\pm$0.02 0.92$\pm$0.02 0.93$\pm$0.01 0.93$\pm$0.01
XGBoost 0.88$\pm$0.04 0.91$\pm$0.01 0.91$\pm$0.01 0.90$\pm$0.01 0.92$\pm$0.01
TabPFN 0.91$\pm$0.02 0.92$\pm$0.02 0.92$\pm$0.02 0.92$\pm$0.01 0.92$\pm$0.02
TabLLM 0.87$\pm$0.06 0.91$\pm$0.01 0.90$\pm$0.01 0.92$\pm$0.01 0.92$\pm$0.01
income CARTE 0.84$\pm$0.09 0.84$\pm$0.02 0.85$\pm$0.03 0.87$\pm$0.01 0.88$\pm$0.01
XGBoost 0.79$\pm$0.03 0.82$\pm$0.02 0.84$\pm$0.01 0.87$\pm$0.01 0.88$\pm$0.00
TabPFN 0.80$\pm$0.04 0.82$\pm$0.04 0.84$\pm$0.01 0.86$\pm$0.01 0.87$\pm$0.01
TabLLM 0.84$\pm$0.01 0.84$\pm$0.02 0.86$\pm$0.01 0.87$\pm$0.00 0.89$\pm$0.01
jungle CARTE 0.71$\pm$0.03 0.80$\pm$0.02 0.81$\pm$0.02 0.86$\pm$0.02 0.90$\pm$0.02
XGBoost 0.78$\pm$0.03 0.81$\pm$0.02 0.84$\pm$0.02 0.87$\pm$0.01 0.91$\pm$0.01
TabPFN 0.78$\pm$0.02 0.81$\pm$0.01 0.84$\pm$0.01 0.88$\pm$0.01 0.91$\pm$0.00
TabLLM 0.71$\pm$0.02 0.78$\pm$0.02 0.81$\pm$0.02 0.84$\pm$0.01 0.89$\pm$0.01
C.2 Detailed Results for TabLLM Datasets
Table 5 shows the dataset specifications and detailed results on performance comparison between CARTE and baselines presented in Hegselmann et al. (2023). The datasets generally contain high fraction of numerical features (four datasets) or categorical columns with low cardinality (eight datasets). In such settings, TabPFN tends to outperform other methods. For the dataset ‘bank’, however, CARTE outperforms other baselines. In particular, the dataset is in line with the 51 datasets that contain both numerical and categorical features with relatively high cardinality of the latter. In a sense, CARTE is in the middle of both TabPFN (suitable for numerical features) and TabLLM (representing information as tokens), with an attentional architecture that has been designed to handle both numerical values and strings.
C.3 Ablation Study on the Components of CARTE
To study the effect of various components of CARTE, we conducted additional experiments in which we exclude or change the associated components. Figure 10 shows the learning curves from the train size of 32 up to 1,024 for each case of excluding or switching the components. For the graph construction with Minhash, we change the feature initialization step with Skrub’s Minhash encoder (Skrub, 2024), which encode string categorical features by applying the MinHash method to n-gram decompositions of strings. The figure shows that each are crucial for gaining the performance of CARTE. In particular, it is interesting to observe the significant decrease with the exclusion of edge information and the attention layer. Since both are essential for leveraging context within a given table, it implies that capturing context is pivotal for attaining the strong performances in predictions. Moreover, the performance gap between CARTE and Minhash confirm that string-level models, that also capture semantic similarity, is important and the use of language models are pivotal for effectively using external information, especially when information given in the table is limited.
(a) Regression
(b) Classification
Figure 10: Ablation on various components of CARTE. The learning curves from the train size of 32 up to 1,024 on various cases excluding or switching the components. Each are crucial for gaining the performance of CARTE. In particular, there is significant decrease with the exclusion of edge information and the attention layer, which are essential for leveraging context within a given table. Moreover, the result with Minhash (A string-level representation without any semantic content, Cerda & Varoquaux, 2022) shows that language models are crucial for effectively using external information.
C.4 The Effect of Oversmoothing
@@
Figure 11 show that representations extracted from deeper layers of the Graph Neural Networks (GNNs) are less useful for prediction on downstream tasks [^46] Here, they are used inside a HistGradientBoosting predictor from scikit-learn.. We interpret this as an effect of oversmoothing, a well known problem in GNNs (Chen et al., 2020a; Rusch et al., 2023).
@@
C.5 Details of the Entity Matching Experiment
@@
C.6 Schema-matching Results on Joint Learning Across Multiple Tables
Figure 12 gives a direct comparison of performance between CARTE with and without schema-matching over 275 different cases in number of source data (ranging from one source to five sources). Each point represents a comparison of the average score over a dataset with different train/test split for a given size of train data. If a point is located below the diagonal line, it indicates a higher performance of x-axis. The figures shows that dots align along the diagonal line, indicating similar performance between CARTE with and without schema-matching (also with p-value of 0.728 for two-sided t-test on difference in means). The results suggest schema-matching is not required for CARTE on transfer across multiple tables.